# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ 5.3. Statistical Baselines
In order to capture privacy-violating (non-LOO) inferences rather than statistical or generalizable (LOO) inferences, the results of the reidentification attack need to be compared to inferences that would be possible in a privacy-preserving counterfactual setting in which the same data are provided, except that a target individual’s record has been removed. In this case, we would compare inferences made about the removed person from the published 2010 Census data to inferences that would be made in the counterfactual world in which the person was removed. Exact comparisons of this sort are difficult because they involve removing a target individual, re-swapping and re-tabulating the data, performing reconstruction and reidentification to make inferences about the individual, and then repeating this process for every person in the United States, or at least a meaningful subset.
@@
Note that the modal guesser is targeting overall accuracy and would perform poorly when guessing the race and ethnicity of minorities within a block. The proportional guesser would perform better with minorities at the expense of overall accuracy, and so both baselines deserve consideration. Comparing the accuracy of reidentification experiments in small, nonmodal populations involving the reconstructed microdata ($\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$) and the statistical baselines (MDG and PRG) estimates the improved inference about these individual due to the used of their actual census responses in the data—the privacy cost of the individual’s participation in the census. In terms of causal inference, the statistical baselines represent the inference accuracy of alternative estimates from the counterfactual world (in which the target person’s data were not used in the publications) whereas the inferences from the reconstructed data represent the inference accuracy of the observed world (in which the target person’s data were used in the publications).
There are two additional points worth making. The first is that the feature set $\{race$, $ethnicity\}$ can also be inferred using the name of the target individual (in which case, the name needs to be part of the observed world and counterfactual inferences). However, since the reconstruction and reidentification algorithms used in this paper do not model this interaction, it is reasonable to omit it from the baseline. What is important for the comparison between the reconstruction and the statistical baseline results is: (1) all else being equal, the estimated privacy cost of being included in the census data publications is the difference in inference accuracy of the reconstruction versus the statistical baseline, [^22] A large difference between the reconstruction accuracy and the statistical baseline accuracy when less information is used, that is, when name is not used for the race/ethnicity inference, is particularly concerning because it directly indicates that confidentiality has already been breached and that additional information from the census response could be precisely revealed in a larger attack based on all SF1 tables. and (2) the contrast between the reconstruction and the statistical baseline inferences gives a lower bound on this privacy cost—the actual confidentiality breach could be larger but not smaller.
@@
We use these statistics to compare our reidentification results with nonstatistical baselines like the HDF itself (i.e., how much better would the reidentification be if the reconstruction were a perfect match to HDF) and with the statistical baselines MDG and PRG. In all cases, the attacker model is the same. The attacker begins with tabular summaries produced from a confidential source, the HDF. Then, the attacker reconstructs record-level images of the persons represented in the tables using either the schema $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$ as illustrated in Table 1. Next, the attacker uses a record-level image that contains identifiable names ($pik$, in this work) and linking variables using the same schema as the reconstructed data, as in the COMCRL and $\text{CEF}_{atkr}$ rows of Table 1, to perform a record-linkage attack that associates the features race and ethnicity with each record in the attacker database. This step produces a record-level database with the features noted in the putative rows of Table 1. Finally, we assess the accuracy of the attack by linking the putative files back to the original confidential data using the features pik and block for linkage, but assessing the accuracy by comparing all features in the row CEF in Table 1.
@@
In summary, the principal indicator of a confidentiality breach is that the precision rate in equation (13) for inferences based on the reconstructed microdata substantially exceeds the precision rate for inferences based on the statistical baselines. The gain in precision, which we will also call the gain in inference accuracy, from the reconstructed data compared to the statistical baseline need not be infinite to constitute a confidentiality breach. The upper-bound on the precision rate of any inference is the precision of that inference based on the CEF itself, since we take it as self-evident that releasing every person record in the CEF for the feature set {block, sex, age, race, ethnicity} is a breach of confidentiality.
6. Reconstruction Solution Variability Is So Limited That Simulation Assessment Is Unnecessary
We assess the agreement of the image of HDF embodied in $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ by comparing those reconstructed microdata files to HDF and CEF themselves using Algorithm 2. However, before assessing the agreement accuracy of $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$, we report results on the solution variability in our reconstruction experiments. If the solutions are highly variable, then simulation assessments like those of [rocher:et:al:2019, dick:etal:2023, ] rather than single-reconstruction assessments, as we propose, would be more appropriate. Table 6 shows every fifth percentile of the cumulative distribution of solvar, which is always assessed using the {block, sex, agebin, race, ethnicity} feature set for the reasons discussed in Section 4.5. The cumulative distribution displayed in the table is over census blocks, not persons. Thus there are $6,207,027/20=310,351$ blocks in each cell of the cumulative distribution, the population of which is shown in the “Population (×$\times 10^{-3}$)” column.
@@ 6. Reconstruction Solution Variability Is So Limited That Simulation Assessment Is Unnecessary
Table 6. Empirical Percentiles for Block-Level Solution Variability
@@
(%) (%) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) $solvar$ (%)
@@
For 70% of all blocks, representing 97,238,000 person records, solution variability is 0. The reconstructed records for those individuals are guaranteed to match their confidential HDF records exactly using the agebin schema. The maximum solution variability for any block in the $75^{th}$ quantile is 3.4%. Since there are 34,272,000 people in this quantile, this means that at most 1,165,000 records in this quantile can differ from their confidential source on race or ethnicity. The maximum cumulative solution variability given any feasible solution for all census blocks, containing all 308,746,000 persons, is just 20.1%. Because this bound applies to any feasible solution, and because the solution variability IP in equation (10) upper-bounds $solvar$ even when using additional constraints, in any other reconstruction solution for $\text{rHDF}_{b,t}$ no more than 20.1% of all records can differ from their HDF record on the value of even a single feature, evaluated on the {block, sex, agebin, race, ethnicity} schema, and all additional age variability in solutions for $\text{rHDF}_{b,t}$ must be within $agebin$ categories. Thus, if agebin is sufficient for high agreement and reconstruction precision, the additional solution variability of the feature age within the feature agebin does not matter.
Finally, we note that [hawes:2022, abowd:hawes:2023, ] report reconstruction agreement and solution variability results based on 32 of the 34 tables shown in Table 2—excluding P8 and P10.252525These results were the basis for presentations the Census Bureau’s Scientific Advisory Committee, whose working group on differential privacy was monitoring the research and implementation decisions underlying the 2020 Census Disclosure Avoidance System. When those two tables are omitted, solution variability is 0 for only 65% of blocks (82.9 million persons), demonstrating empirically the mathematical property of the solution variability IP (10)—as more tables are added the solutions become less variable. In this case, two tables with additional {race, ethnicity} data for adults increased the number of 0 solution variability records by 14.3 million. That is 14.3 million additional people for whom the attacker is certain that the reconstructed record matches the HDF record exactly using the agebin schema.
@@
Size Blocks (×$\times 10^{-3}$) (%) Blocks (×$\times 10^{-3}$) (%) (%)
@@
Table 7 shows the percent of population uniques by census block size, for all blocks, and for those with 0 reconstruction solution variability. Over 99% of the population in the smallest blocks are in blocks where there is only one set of microdata records consistent with the published tables, meaning that a population unique on {block, sex, agebin} is identified with certainty. Even among blocks with population from 10 to 49, 71.6% of individuals are in blocks with unique reconstructed microdata, meaning that in those blocks a population unique on {block, sex, agebin} is also identified with certainty. Among blocks with 50-99 individuals, 0 solution variability blocks are less common, but still substantial. Overall, 70% of all blocks had no solution variability, representing 97,238,000 persons, as shown in Table 6. Since a record that is population unique on the feature set{block, sex, agebin} must also be unique on {block, sex, age}, no access to confidential data is required to learn the uniqueness of the most basic set of quasi-identifiers—residence location, sex, and age—and, therefore, the associated feature values {race, ethnicity} in the confidential HDF. Traditional SDL considers this exposure of population uniqueness in released data de facto prohibited disclosure: “[i]f a unit is population unique then disclosure will occur if a snooper knows it is unique” [duncan:etal:2011:statistical], pp. 42-43). Swapping the confidential CEF to create the confidential HDF introduces some uncertainty in the inference about the learned feature values {race, ethnicity}, but that uncertainty is upper bounded by the swap rate, which the Office of National Statistics in the United Kingdom has acknowledged by publishing the swap rate and stating that adding additional noise to every tabular summary was also required, effectively mimicking the consequences of the U.S. Census Bureau’s 2020 DAS without using a differential privacy framework [ONS:2023, dove:et:al:ONSApplying].
7. Reconstructed Data Strongly Agree with the Confidential Data
@@
Table 8 shows the agreement match results for all census blocks (top panel) and for each of the seven census block-size intervals we considered throughout this work. Note that since HDF-HDF and CEF-CEF comparisons always yield 100% agreement, these rows are not shown. Consider first the HDF-CEF comparison. We find that the error introduced by the 2010 SDL methodology is relatively small. Less than $4\%$ of all person records disagree using either the age or agebin schema. The reconstruction $\text{rHDF}_{b,t}$ agrees with the HDF using exact age for 48.5% of records, and using binned age for 95.2% of records. Comparing $\text{rHDF}_{b,t}$ to CEF decreases the agreement rates only slightly for binned age (from 95.2% to 91.9%) or exact age match (from 48.5% to 46.6%), indicating that many of the reconstructed ages are off by a small number of years, but that this error in the reconstruction is entirely within the narrow bands of the agebin schema, implying that reidentification based on binned age will be almost as reliable as using exact age, especially for the population 21 and under, for which the agebin schema is the same as the exact age schema.
The $\text{rHDF}_{b}$ reconstruction agreement rates with HDF and CEF are all slightly less than the respective $\text{rHDF}_{b,t}$ agreement rates. Thus, using tract-level information improves the reconstruction accuracy relative to using only block-level tables even when using the binned-age feature. The agreement gain of $\text{rHDF}_{b,t}$ over $\text{rHDF}_{b}$ with respect to the HDF of 2.3 percentage points when comparing with the binned-age feature ($=95.2-92.9$) is entirely an improvement in {race, ethnicity} matching because there is never any solution variability in the features {block, sex, agebin}—we always find the same the solution for these features once the tables P12 and P14 in Panel A of Table 2 are used because these two tables fully saturate this feature set. The agreement gain with respect to the HDF of 4.3 percentage points ($=48.5-44.2$) can, therefore, be decomposed into 2.3 percentage points of improved agreement on {race, ethnicity} and 2.0 percentage points of improved agreement on age due to the information in the tract-level tables of shown in Panel B Table 2.
@@
Data ($L$-$R$ in Block Population Agreement (×$\times 10^{-3}$) Agreement (%)
Algorithm 2) Size (×$\times 10^{-3}$) Exact Age Binned Age Exact Age Binned Age
@@
Next consider how agreement of $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ with HDF and CEF varies as a function of census block size (Table 8). Because low-population blocks are much more likely to have had the targeted 2010 Census SDL treatment applied, agreement between the HDF and CEF is only 74.0% in blocks with populations of 1-9 persons. This does not mean the SDL was effective; as we show in Section 8, when a record linkage is made, additional attributes about a record can be accurately learned. Furthermore, we made no attempt to undo household record swaps. Using the binned-age schema, $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ also agree with the CEF for 74.0% of records in the 1-9 person block size, as the next two rows of Table 8 show because virtually all persons who live in these low-population blocks have 0 solution variability (see Table 7). For persons living in blocks with population of 10-49, the HDF agrees with the CEF for 94.1% using the binned-age schema. $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ agree with the CEF for 93.0% and 92.2%, respectively. However, once again, because the vast majority of these persons have 0 solution variability, the reconstructions agree with the HDF 98.8% and 97.8%, respectively. Thus, the certainty with which low-population blocks can be reconstructed—there is almost always only one solution to the reconstruction and that one solution is provably identical to the HDF record— results in extremely high empirical agreement between the reconstructed and confidential HDFs.
@@
We have two final comments on these reconstruction and solution variability results. First, both reconstruction and solution variability can be computed using only public data, meaning that an attacker would be able to use these methods to target particular blocks for which the reconstruction is unique—access to confidential data is not required to confirm that these records are exact images of the HDF. Second, inferences about persons in low-population blocks tend to be more sensitive to the absence of a record because leaving out a unique or near-unique record type in such blocks is much more likely to change the inference than leaving out one of a large number of records of the same type, say the modal {race, ethnicity}. When there is 0 solution variability, as there is in the vast majority of blocks with fewer than 50 people, this effect is even stronger because in these blocks a record unique in $\text{rHDF}_{b}$ is provably unique in the HDF, while in blocks with greater solution variability, this may not be the case.
Taken together, the patterns in Table 8 imply that the $\text{rHDF}_{b,t}$ matches the CEF as well as the actual HDF matches the CEF for smaller blocks, which means that the Census Bureau was correct in considering low-population geographies a significant disclosure risk in microdata but incorrect in assuming that tabular aggregation could mitigate this risk. Thus, Table 8 demonstrates conclusively that the 2010 Census disclosure avoidance system was fatally flawed by the failure to use SDL treatments for tabular data that were consistent with rules for microdata. [^27] In other words, continuing with the 2010 technology, instead of designing a new disclosure avoidance system for 2020, would likely have involved both dramatically increasing the swap rate and heavy use of suppression of block level tables. Sampling, aggregating geographic identifiers to populations of 100,000+, and aggregating demographic groups to national populations of 10,000+ probably protected the 2010 Public-Use Microdata Sample (PUMS) adequately considering that, unlike the 2000 Census PUMS, there were no long-form data to include. As we show in the next section, these reconstructed microdata permit prohibited confidentiality breaches just as the 2010 PUMS would have permitted had the Census Bureau used only its 2010 Census tabular data SDL treatment for the 2010 PUMS.

### v2

@@ 5.3. Statistical baselines
In order to capture confidentiality-violating (non-LOO) inferences rather than statistical or generalizable (LOO) inferences, the results of the reidentification attack need to be compared to inferences that would be possible in a privacy-preserving counterfactual setting in which the same data are provided, except that a target individual’s record has been removed. In this case, we would compare inferences made about the removed person from the published 2010 Census data to inferences that would be made in the counterfactual world in which the person was removed. Exact comparisons of this sort are difficult because they involve removing a target individual, re-swapping and re-tabulating the data, performing reconstruction and reidentification to make inferences about the individual, and then repeating this process for every person in the United States, or at least a meaningful subset.
@@
Note that the modal guesser is targeting overall accuracy and would perform poorly when guessing the race and ethnicity of minorities within a block. The proportional guesser would perform better with minorities at the expense of overall accuracy, and so both baselines deserve consideration. Comparing the accuracy of reidentification experiments in small, nonmodal populations involving the reconstructed microdata ($\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$) and the statistical baselines (MDG and PRG) estimates the improved inference about these individuals due to the used of their actual census responses in the data—the privacy cost of the individual’s participation in the census. In terms of causal inference, the statistical baselines represent the inference accuracy of alternative estimates from the counterfactual world (in which the target person’s data were not used in the publications) whereas the inferences from the reconstructed data represent the inference accuracy of the observed world (in which the target person’s data were used in the publications).
There are two additional points worth making. The first is that the feature set $\{race$, $ethnicity\}$ can also be inferred using the name of the target individual (in which case, the name needs to be part of the observed world and counterfactual inferences). However, since the reconstruction and reidentification algorithms used in this paper do not model this interaction, it is reasonable to omit it from the baseline. What is important for the comparison between the reconstruction and the statistical baseline results is: (1) all else being equal, the estimated privacy cost of being included in the census data publications is the difference in inference accuracy of the reconstruction versus the statistical baseline, [^24] A large difference between the reconstruction accuracy and the statistical baseline accuracy when less information is used, that is, when name is not used for the race and ethnicity inference, is particularly concerning because it directly indicates that confidentiality has already been breached and that additional information from the census response could be precisely revealed in a larger attack based on all SF1 tables. and (2) the contrast between the reconstruction and the statistical baseline inferences gives a lower bound on this privacy cost—the actual confidentiality breach could be larger but not smaller.
@@
We use these statistics to compare our reidentification results with nonstatistical baselines like the HDF itself (i.e., how much better would the reidentification be if the reconstruction were a perfect match to HDF) and with the statistical baselines MDG and PRG. In all cases, the attacker model is the same. The attacker begins with tabular summaries produced from a confidential source, the HDF. Then, the attacker reconstructs record-level images of the persons represented in the tables using either the schema $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$ as illustrated in Table 1. Next, the attacker uses a record-level image that contains identifiable names ($pik$, in this work) and linking variables using the same schema as the reconstructed data, as in the COMRCL and $\text{CEF}_{atkr}$ rows of Table 1, to perform a record-linkage attack that associates the features race and ethnicity with each record in the attacker database. This step produces a record-level database with the features noted in the putative rows of Table 1. Finally, we assess the accuracy of the attack by linking the putative files back to the original confidential data using the features pik and block for linkage, but assessing the accuracy by comparing all features in the row CEF in Table 1.
@@
In summary, the principal indicator of a confidentiality breach is that the precision rate in Equation 13 for inferences based on the reconstructed microdata substantially exceeds the precision rate for inferences based on the statistical baselines. The gain in precision, which we will also call the gain in inference accuracy, from the reconstructed data compared to the statistical baseline need not be infinite to constitute a confidentiality breach. The upper-bound on the precision rate of any inference is the precision of that inference based on the CEF itself, since we take it as self-evident that releasing every person record in the CEF for the feature set {block, sex, age, race, ethnicity} is a breach of confidentiality.
6. Reconstruction Solution Variability Is So Limited
That Simulation Assessment Is Unnecessary
6.1. Solution variability
In this section we show that the reconstructed 2010 Census microdata have remarkably little solution variability. That is, about one-third of the reconstructions are exact for the census block-level variables (binned-age schema). The remaining records have very little error—none at all in census block, sex, and binned age, and a small amount in race and ethnicity. Given these results, there is no real advantage to generating additional reconstruction solutions. The main conclusions can be drawn from the 97 million records that are solved exactly. Hence, given the very low solution variability, we assess the agreement of the full image of HDF embodied in $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ by comparing our reconstructed microdata directly to the confidential HDF and CEF using Algorithm 2. Because the solutions are not highly variable, using a single reconstruction is computationally simpler than simulation assessments like those of rocher:et:al:2019; dick:etal:2023 while still allowing the quantification of worst-case variability.
Table 6 shows every fifth percentile of the cumulative distribution of solvar, which is always assessed using the {block, sex, agebin, race, ethnicity} feature set. The cumulative distribution shown in the table is computed over census blocks, not persons. Thus there are $6,207,027/20=310,351$ blocks in each cell of the cumulative distribution, the population of which is shown in the “Population (×$\times 10^{3}$)” column.
@@ 6.1. Solution variability
Table 6. Empirical Percentiles for Census Block-Level Solution Variability
@@
(%) (%) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) $solvar$ (%)
@@
For 70% of all blocks, representing 97,238,000 person records, solution variability is 0. The reconstructed records for those individuals are guaranteed to match their confidential HDF records exactly using the agebin schema. The maximum solution variability for any block in the $75^{th}$ quantile is 3.4%. Since there are 34,272,000 people in this quantile, this means that at most 1,165,000 records in this quantile can differ from their confidential source on race or ethnicity. The maximum cumulative solution variability given any feasible solution for all census blocks, containing all 308,746,000 persons, is just 20.1%. Because this bound applies to any feasible solution, and because the solution variability IP in Equation 10 upper-bounds $solvar$ even when using additional constraints, in any other reconstruction solution for $\text{rHDF}_{b,t}$ no more than 20.1% of all records can differ from their HDF record on the value of even a single feature, evaluated on the {block, sex, agebin, race, ethnicity} schema, and all additional age variability in solutions for $\text{rHDF}_{b,t}$ must be within $agebin$ categories. Thus, if agebin is sufficient for both high agreement rates and high reidentification precision, the additional solution variability of the feature age within the feature agebin does not matter for our assessments.
Finally, we note that hawes:2022; abowd:hawes:2023 report reconstruction agreement and solution variability results based on 32 of the 34 tables shown in Table 2—excluding P8 and P10.272727These results were the basis for presentations the Census Bureau’s Scientific Advisory Committee, whose working group on differential privacy was monitoring the research and implementation decisions underlying the 2020 Census Disclosure Avoidance System. When those two tables are omitted, solution variability is zero for only 65% of blocks (82.9 million persons), demonstrating empirically the mathematical property of the solution variability IP 10—as more tables are added the solutions become less variable. In this case, two tables with additional race data for all persons and persons age 18 and older increased the number of zero solution variability records by 14.3 million. That is 14.3 million additional people for whom the attacker is certain that the reconstructed record matches the HDF record exactly using the agebin schema.
Our solution variability results imply that for much of the U.S. population the record-level image of the features used to create the census tract and block-level data shown in Table 2 is essentially an exact copy of HDF on the binned-age schema. Thus, the SF1 data shown in Table 2 are equivalent to the microdata HDF records for the {block, sex, agebin, race, ethnicity} feature set. The release of these microdata was prohibited by the 2010 Census disclosure avoidance rules (mckenna:2019:microdata).
Permission to release the reconstructed HDF was approved in 2022 under clearance number CBDRB-FY22-DSEP-004. The public replication archive for this paper includes a sample of the reconstructed records in 29 tracts with zero solution variability, all necessary inputs and programs to reconstruct the entire 308,745,538 person records for $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ from public data, and everything required to confirm our solution variability results.
@@
Size Blocks (×$\times 10^{3}$) (%) Blocks (×$\times 10^{3}$) (%) (%)
@@
Table 7 shows the percent of population uniques by census block size, for all blocks, and for those with zero reconstruction solution variability. Over 99% of the population in the smallest blocks are in blocks where there is only one set of microdata records consistent with the published tables, meaning that a population unique on {block, sex, agebin} is identified with certainty. Even among blocks with population from 10 to 49, 71.6% of individuals are in blocks with unique reconstructed microdata, meaning that in those blocks a population unique on {block, sex, agebin} is also identified with certainty. Among blocks with 50-99 individuals, zero solution variability blocks are less common, but still substantial. Overall, 70% of all blocks had no solution variability, representing 97,238,000 persons, as shown in Table 6. Since a record that is population unique on the feature set{block, sex, agebin} must also be unique on {block, sex, age}, no access to confidential data is required to learn the uniqueness of the most basic set of quasi-identifiers—residence location, sex, and age—and, therefore, the associated feature values {race, ethnicity} in the confidential HDF. Traditional SDL considers this exposure of population uniqueness in released data de facto prohibited disclosure: “[i]f a unit is population unique then disclosure will occur if a snooper knows it is unique” duncan:etal:2011:statistical, pp. 42-43). Swapping the confidential CEF to create the confidential HDF introduces some uncertainty in the inference about the learned feature values {race, ethnicity}, but that uncertainty is upper bounded by the swap rate, which the Office of National Statistics in the United Kingdom has acknowledged by publishing the swap rate and stating that adding additional noise to every tabular summary was also required, effectively mimicking the consequences of the U.S. Census Bureau’s 2020 DAS without using a differential privacy framework (ONS:2023; ONS:2023:appendix; spicer:2020).
6.2. Reconstructed data agreement with the confidential data
@@
Table LABEL:tab:reconstruction_by_blocksize shows the agreement match results for all census blocks (top panel) and for each of the seven census block-size intervals we considered throughout this work. Note that since HDF-HDF and CEF-CEF comparisons always yield 100% agreement, these rows are not shown. Consider first the HDF-CEF comparison. We find that the error introduced by the 2010 SDL methodology is relatively small. Less than $4\%$ of all person records disagree using either the age or agebin schema. The reconstruction $\text{rHDF}_{b,t}$ agrees with the HDF using exact age for 48.5% of records, and using binned age for 95.2% of records. Comparing $\text{rHDF}_{b,t}$ to CEF decreases the agreement rates only slightly for binned age (from 95.2% to 91.9%) or exact age match (from 48.5% to 46.6%), indicating that many of the reconstructed ages are off by a small number of years, but that this error in the reconstruction is entirely within the narrow bands of the agebin schema, implying that reidentification based on binned age will be almost as reliable as using exact age, especially for the population 21 and under, for which the agebin schema is the same as the exact age schema.
The $\text{rHDF}_{b}$ reconstruction agreement rates with HDF and CEF are all slightly less than the respective $\text{rHDF}_{b,t}$ agreement rates. Thus, using tract-level information improves the reconstruction accuracy relative to using only block-level tables even when using the binned-age feature. The agreement gain of $\text{rHDF}_{b,t}$ over $\text{rHDF}_{b}$ with respect to the HDF of 2.3 percentage points when comparing with the binned-age feature ($=95.2-92.9$) is entirely an improvement in {race, ethnicity} matching because there is never any solution variability in the features {block, sex, agebin}. We always find the same the solution for these features once the tables P12 and P14 in Panel A of Table 2 are used because these two tables fully saturate this feature set. The agreement gain with respect to the HDF of 4.3 percentage points ($=48.5-44.2$) can, therefore, be decomposed into 2.3 percentage points of improved agreement on {race, ethnicity} and 2.0 percentage points of improved agreement on age due to the information in the tract-level tables of shown in Panel B Table 2.
@@
Data ($L$-$R$ in Block Population Agreement (×$\times 10^{3}$) Agreement (%)
Algorithm 2) Size (×$\times 10^{3}$) Exact Age Binned Age Exact Age Binned Age
@@
Next consider Table LABEL:tab:reconstruction_by_blocksize, which illustrates how agreement of $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ with HDF and CEF varies as a function of census block size. Because low-population blocks are much more likely to have had the targeted 2010 Census SDL treatment applied, agreement between the HDF and CEF is only 74.0% in blocks with populations of 1-9 persons. This does not mean the SDL was effective; as we show in Section 7, when a record linkage is made, additional attributes about a record can be accurately learned. Furthermore, we made no attempt to undo household record swaps. Using the binned-age schema, $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ also agree with the CEF for 74.0% of records in the 1-9 person block size, as the next two rows of Table LABEL:tab:reconstruction_by_blocksize show because virtually all persons who live in these low-population blocks have zero solution variability (see Table 7). For persons living in blocks with population of 10-49, the HDF agrees with the CEF for 94.1% using the binned-age schema. $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ agree with the CEF for 93.0% and 92.2%, respectively. However, once again, because the vast majority of these persons have zero solution variability, the reconstructions agree with the HDF 98.8% and 97.8%, respectively. Thus, the certainty with which low-population blocks can be reconstructed—there is almost always only one solution to the reconstruction and that one solution is provably identical to the HDF record— results in extremely high empirical agreement between the reconstructed and confidential HDFs.
@@
We have two final comments on these reconstruction and solution variability results. First, both reconstruction and solution variability can be computed using only public data, meaning that an attacker would be able to use these methods to target particular blocks for which the reconstruction is unique—access to confidential data is not required to confirm that these records are exact images of the HDF. Second, inferences about persons in low-population blocks tend to be more sensitive to the absence of a record because leaving out a unique or near-unique record type in such blocks is much more likely to change the inference than leaving out one of a large number of records of the same type, say the modal {race, ethnicity}. When there is zero solution variability, as there is in the vast majority of blocks with fewer than 50 people, this effect is even stronger because in these blocks a record unique in $\text{rHDF}_{b}$ is provably unique in the HDF, while in blocks with greater solution variability, this may not be the case.
Taken together, the patterns in Table LABEL:tab:reconstruction_by_blocksize imply that the $\text{rHDF}_{b,t}$ matches the CEF as well as the actual HDF matches the CEF for smaller blocks, which means that the Census Bureau was correct in considering low-population geographies a significant disclosure risk in microdata but incorrect in assuming that tabular aggregation could mitigate this risk. Thus, Table LABEL:tab:reconstruction_by_blocksize demonstrates conclusively that the 2010 Census disclosure avoidance system was fatally flawed by the failure to use SDL treatments for tabular data that were consistent with rules for microdata. [^29] In other words, continuing with the 2010 technology, instead of designing a new disclosure avoidance system for 2020, would likely have involved both dramatically increasing the swap rate and heavy use of suppression of block level tables. For microdata, and specifically the Public-Use Microdata Sample (PUMS), the requirements were very different: sampling, aggregating geographic identifiers to populations of 100,000+, and aggregating demographic groups to national populations of 10,000+ probably protected the PUMS adequately considering that, unlike the 2000 Census PUMS, there were no long-form data to include. As we show in the next section, these reconstructed microdata permit prohibited confidentiality breaches just as the 2010 PUMS would have permitted had the Census Bureau used only its 2010 Census tabular data SDL treatment for the 2010 PUMS.